# Plasma acceleration of polarized particle beams

### v1

@@ 3.2 Positrons
The work on polarized positron beams is currently divided between the production and the acceleration of these beams. In contrast to electrons, no pre-polarized targets of positrons can be produced, therefore production via processes like Breit-Wheeler pair creation is necessary.
@@
An alternative method is described by Li et al.: when a high-intensity CP laser pulses interacts with a longitudinally polarized, relativistic electron beam, a highly polarized positron beam can be created via the non-linear Breit-Wheeler process [59]. The polarization is transferred from the electrons to the positrons by high-energy photons. The method is shown to produce positron beams with 40%-60% polarization and low angular divergence on a femtosecond time scale.
Liu et al. found the polarization of intermediate gamma quanta to be of significance for the polarization of electron-positron pairs produced in the non-linear Breit-Wheeler process [60]. The setup is shown in Fig. 6. They find that an average polarization of 30% can be obtained, which is then accelerated in a wakefield driven by a hollow electron beam without significant depolarization.
Figure 6: (a) The hollow driver and an electron seed beam propagate in $z$-direction. (b) The driver beam generates the wakefield structure, and the seed is irradiated by a two-color laser pulse in order to produce positrons. (c) The polarized positrons are then trapped and accelerated in the wakefield. Reproduced under the terms of the CC-BY license from [60]. Copyright 2022, The Authors, published by American Physical Society.
A method utilizing laser-solid interaction has been proposed by Xue et al. [61]. A high-intensity laser pulse is used to ionize a solid foil. The accelerated electrons are accelerated and generate gamma quanta via non-linear Compton scattering which decay into polarized positrons in a quasi-static magnetic field. The angle of irradiation is found to determine the configuration of the quasi-static field, which in turn, allows tuning of the polarization degree. The scheme delivers polarization on a 70%-level with 0.1 nC per shot. Zhao et al. investigated cascaded Compton scattering and Breit-Wheeler pair production on a more general level with respect to polarization transfer in [62].
@@ 4.1 Magnetic Vortex Acceleration
In the context of polarized protons, MVA was first investigated by Jin et al. in [35] by means of particle-in-cell simulations. They investigated the interaction of a multi-PW laser pulse in the range $a_{0}=25-100$ with a pre-polarized HCl target. The target has an electron density of $0.36n_{\mathrm{crit}}$ and is has a flat-top profile with steep edges. In the case of $a_{0}=25$, 0.26 nC of protons with up to 53 MeV and a polarization of 82% were observed. An increase in laser energy was shown to reduce the polarization: for a laser pulse with $a_{0}=100$, 3.1 nC were accelerated to 152 MeV, however with a lower polarization of only 56%.
A publication by Reichwein et al. studied the effect of density-down ramps at the end of the target on the final beam quality [67]. The same HCl target was used, however with different lengths of down-ramps in the range of 0-80 $\mu$m. The polarization of the final proton beam was shown to remain largely robust against changes of the ramp length. Longer ramps mostly affected the spatial focusing of the beam as the transverse extent of the plasma channel and, in turn, the induced fields are changed. Ramps of significant length can, however, lead to reduced beam polarization, since a longer ramp changes the focusing of protons into the beam, introducing also particles of different spin orientation.
@@
In the space between the two plasma channels, an accelerating region with weaker depolarizing fields is formed (cmp. current density in Fig. 7). This leads to the formation of a central filament which obtains energies similar to the comparable single-pulse MVA setup, but with higher polarization. In [68], polarization of 77% was obtained for two pulses with $a_{0}=100$, while a single-pulse setup of the same total laser energy only led to 64%. The dual-pulse scheme was also shown to improve on the angular spectrum of the accelerated beam, as instabilities were mitigated by the presence of the second laser pulse.
Figure 7: Transverse cut of the current density profile of the dual-pulse MVA scheme. Note the two distinct plasma channels with an accelerating region in their middle. Reproduced under the terms of the CC-BY license from [68]. Copyright 2022, The Authors, published by American Physical Society.
@@ 4.2 Collisionless Shock Acceleration
In [70], Yan and Ji first suggested the use of CSA for polarized proton beams. Their setup consisted of a laser pulse of $a_{0}=20-80$ irradiating a 2 $\mu$m Carbon foil which is placed in front of a near-critical, polarized HCl target. Proton beams with tens of MeV and a polarization on a 90%-level were obtained. A subsequent study by Yan et al. investigated the use of a microstructured foil [71], where the structures were utilized to increase the temperature and the number of hot electrons. Accordingly, the energy of the reflected protons was increased.
The aforementioned studies have been extended by Reichwein et al. to the case of polarized 3He and to higher laser intensities [72]. A schematic of the setup is shown in Fig. 8. Similarly to the results with protons, high polarization is achieved with CSA. In particular, the degree polarization remains in the range of 90% for intensities of $a_{0}=$ 100-200. Moreover, the publication investigated the influence of radiation reaction on the acceleration process. As the electrons radiate part of their energy, the electromagnetic field structure of the system and, subsequently, the ion motion is changed. This leads to a reduction in energy, but an increase in beam charge, since more Helium ions can be trapped inside the accelerating field. The polarization remains largely unaffected by radiation reaction.
Figure 8: Schematic of the setup for Collisionless Shock Acceleration. A solid Carbon foil is placed in front of the near-critical density, polarized Helium-3 target. Reproduced under the terms of the CC-BY license from [72]. Copyright 2024, The Authors, published by IOP Publishing Ltd.
An alternative pathway to CSA was suggested by Yan et al. in [73], where no additional foil is necessary to drive CSA. A CO2 laser with a wavelength of 10 $\mu$m was used, which gives a critical plasma density comparable to the density achieved in pre-polarized targets created by photo-dissociation. This allows a laser pulse in the intensity range of $10^{17}-10^{18}$ W/cm2 to be used. The shock front is created by the long laser pulse (3 ps) that piles up protons. A spectral peak energy of approximately 2 MeV was found, with a cut-off energy of about 7 MeV. The polarization remained $>80\%$ in a large parameter range.
@@ 4.3 Other mechanisms
Hützen et al. also considered the acceleration of ions using a bubble-channel structure for multi-PW laser systems [74]. The acceleration scheme is based on work by Shen et al. [75]. Both HCl and HT plasma with a densities in the range of $10^{19}$-$10^{21}$ cm-3 and a length of 600 $\mu$m are considered. The plasma is irradiated by a high-intensity laser pulse of up to $a_{0}=223$, leading to the formation of a wide plasma channel. At the leading edge of this channel structure, polarized particles are piled up and accelerated to energies in the multi-GeV range. The investigation of this mechanism was extended by Li et al. in [76]. Here the protons are pre-accelerated by a CP laser pulse, and subsequently trapped and further accelerated in the front region of a wakefield bubble. The ratio of hydrogen and tritium is found to be crucial for the efficiency of the acceleration. If chosen correctly, a polarized proton beam can be accelerated up to GeV-level in the wakefield structure.
The acceleration of polarized ions from hypothetical, overdense targets was considered by Hützen et al. in [36]. Two sets of PIC simulations were carried out: at first a fully polarized Hydrogen layer of 1 $\mu$m thickness and $128n_{\mathrm{crit}}$ density was considered (see Fig. 9). The second set of simulations considered a 2.5 $\mu$m aluminium layer at $35n_{\mathrm{crit}}$ with a shorter polarized proton layer with 0.5 $\mu$m thickness. The dominant acceleration mechanism in both setups was found to be Target Normal Sheath Acceleration (TNSA). In the former case, preservation of proton polarization for up to 0.24 ps was observed, while for the latter case it was found to be preserved for 0.18 ps seconds.
Figure 9: Proton spin polarization for the pure Hydrogen foil (a)-(b), as well as the Aluminium-Hydrogen composite target (c)-(d). Reproduced under the terms of the CC-BY license from [36]. Copyright 2019, The Authors, published by Cambridge University Press.
@@ 5 Polarized gamma quanta
In this section, we discuss the generation of polarized gamma-ray photons in the interaction between a high-intensity laser pulse and plasmas (or charged particle beams). The interest in studying polarized gamma-ray photons driven by high-intensity lasers is motivated by the fact that both the practical application of the polarized quanta and the potential multi-messenger laser-driven plasmas. The polarized gamma-ray photons have practical applications including nuclear physics [77], high-energy physics [78, 79], and astrophysics [80, 81]. Moreover, the photon polarization can carry the additional degree of freedom of information to help retrieve the in-situ electron dynamics in the ultrarelativistic laser or beam driven plasmas [82].
@@ 5.1 Polarized Multi-GeV gamma-photon beams via Single-Shot Laser-Electron Interaction
Recently, by considering the photon polarization effect, Li et al. [83] investigated the generation of circularly polarized (CP) and linearly polarized (LP) $\gamma$-rays via the interaction of an ultraintense laser pulse with a spin-polarized counter-propagating ultrarelativistic electron beam. Pre-polarized energetic electrons collide with the laser pulse to experience nonlinear Compton scattering in the quantum radiation-dominated regime as shown in Fig. 10, where a GeV pre-polarized electron beam is set to collide with a counter-propagating laser pulse with a relativistic intensity. The study develops a Monte Carlo method based on the locally constant field approximation (LCFA) that employs electron-spin-resolved probabilities for polarized photon emissions. Li et al. [83] show efficient ways for the transfer of electron polarization to high-energy photon polarization. The results demonstrate that multi-GeV CP (LP) $\gamma$-rays with polarization of up to about 95% can be generated by a longitudinally (transversely) spin-polarized electron beam, with a photon flux meeting the requirements of recent proposals for the vacuum birefringence measurement in ultra-strong laser fields. The high-energy, high-brilliance, and high-polarization $\gamma$-rays generated in this process are also beneficial for other applications in high-energy particle physics, light sources, and laboratory astrophysics. Recently, using the quantum operator method by Baier and Katkov to calculate the polarization-resolved probabilities within the quasi-classical approach and LCFA, Chen et al. [84] presented a detailed derivation of fully polarization-resolved probabilities for high-energy photon emission and $e^{-}e^{+}$ pair production in ultra-strong laser fields. These probabilities, accounting for both electron spin and photon polarization of incoming and outgoing particles, are crucial for the development of QED Monte Carlo simulations and QED-particle-in-cell codes.
@@
Figure 10: Scenarios of generating CP and LP $\gamma$-rays via nonlinear Compton scattering. (a) An arbitrarily polarized (AP) laser pulse propagating along $+z$ direction and head-on colliding with a longitudinally spin-polarized electron bunch produces CP $\gamma$-rays rays. (b) An elliptically polarized laser pulse propagating along $+z$ direction and colliding with a transversely spin-polarized (TSP) electron bunch produces LP $\gamma$-rays. The major axis of the polarization ellipse is along the $x$ axis. Reproduced with permission from [83]. Copyright 2020, American Physical Society.
@@
In subsequent studies, Wan et al. [85] investigated the interaction between an unpolarized 10 GeV electron beam and a counter-propagating linearly polarized laser pulse (with a peak intensity $a_{0}\approx 50$ and pulse duration $\tau\approx 50\,$fs) using the same semiclassical Monte Carlo approach. They found that high-energy, linearly polarized $\gamma$-ray photons are generated during the interaction through nonlinear Compton scattering, with an average polarization degree exceeding 50%. These photons further interact with the laser fields, leading to electron-positron pair production via the nonlinear Breit-Wheeler process. Photon polarization is found to enhance the pair production yield by over 10%. To understand the influence of photon polarization on pair production, Dai et al. [86] using fully polarization-resolved Monte Carlo simulations to examined the correlation between photon and electron (positron) polarization in the multiphoton Breit-Wheeler process, where a 10 GeV electron beam is set to collide with a counter-propagating elliptically polarized laser pulse with a peak intensity $a_{0}=100$, pulse duration $\tau\approx 27\,$fs, and the ellipticity $\epsilon=0.03$. They found that the polarization of produced $e^{-}e^{+}$ is reduced by 35% and the pair yield is decreased by 13% if the intermediate photon polarization is taken into account.
@@ 5.2 Polarized high-energy brilliant gamma-ray sources via laser-plasma interaction
Figure 11: Scenario for the generation of LP $\gamma$-ray photons via nonlinear Compton scattering. An ultra-strong LP laser pulse, polarized along the $y$-axis and propagating along the $x$-direction, irradiates a near-critical density hydrogen plasma followed by an ultra-thin Al foil. Reproduced under the terms of the CC-BY license from [87]. Copyright 2020, The Authors, published by AIP Publishing.
@@
It is worth emphasizing that the numerical study of polarized $\gamma$-ray photons via single-shot laser-electron interaction is based on a test particle simulation [83, 84, 85, 86]. This means that collective effects of charged particles, such as self-generated electric and magnetic fields, are not taken into account. Recently, Xue et al. [87] studied about the generation of highly-polarized high-energy brilliant $\gamma$-rays via laser-plasma interaction, where plasma collective effects are self-consistently considered by implementing the electron-spin-resolved probabilities for polarized photon emission in PIC simulations. By utilizing PIC simulations, Xue et al. [87] show that when an ultraintense LP laser pulse irradiates a near-critical-density (NCD) plasma target followed by an ultra-thin aluminum (Al) foil, the electrons in the NCD plasma are first accelerated by the driving laser via direct laser acceleration to ultrarelativistic energies and then collide with the light pulse reflected by the Al foil. The PIC simulations of this all-optical scheme show that brilliant LP $\gamma$-ray photons are produced via nonlinear Compton backscattering with photon energy up to hundreds of MeV and an average LP polarization degree of $\sim 70\%$. This high LP polarization degree is reasonable since the driving laser pulse is a fully polarized light with a LP degree of $100\%$. Alternatively, instead of using the double layer target in Fig. 11 to generate $\gamma$-ray photons via nonlinear Compton backscattering, Xue et al. [87] also calculate photon emission by considering a conical gold target filled with an NCD hydrogen plasma. When the ultraintense LP light impinges the NCD target inside the Au cone, the bulk plasma electrons are pushed forward and the consequent current density sustains a strong quasi-static self-generated magnetic field. For the forward moving electrons, the self-generated magnetic field provides a restoring force to enable electrons’ betatron oscillation inside guided plasma channel in the conical Au target, in which emitted $\gamma$-ray photons concentrate in the angular region $\theta<20^{\circ}$, and the cutoff energy exceeds 450 MeV. In this scheme, the LP degree can reach $\sim 50\%$ for the energetic part of emitted photons.
By employing the similar scheme of the combination of laser wakefield acceleration and plasma mirror as shown in Fig. 12, Wang et al. [88] found that a brilliant CP $\gamma$-ray beam could be generated in a weakly nonlinear Compton scattering regime with moderate laser intensities of $a_{0}\approx 3$, where the helicity of the driving laser pulse is transferred to the emitted $\gamma$-photons. The authors claimed that the calculated average polarization degree of $\gamma$-photons reaches $\sim 61\%(20\%)$ with a peak brilliance of $\gtrsim 10^{21}$ photons/($\cdot$ mm2$\cdot$ mrad2$\cdot$ 0.1% BW) around 1 MeV (100 MeV) [88]. Besides, Wang et al. also investigated the transfer mechanism of spin angular momentum during the transition from linear to nonlinear processes [89]. Their findings suggest that to achieve high-energy, high-brilliance, and high-polarization gamma rays, increasing the laser intensity is crucial for an initially spin-polarized electron beam. However, for an initially unpolarized electron beam, it is also necessary to increase the electron beam’s energy, in addition to boosting laser intensity.
@@
5.3 Application: deciphering in situ electron dynamics of ultrarelativistic plasma
From Li et al. [83] and Xue et al. [87], one can find that $\gamma$-ray photons with high LP degree of $P_{\mathrm{LP}}>70\%$ can be produced via nonlinear Compton backscattering. These highly polarized $\gamma$-ray photons have various applications in laboratory astrophysics and high-energy physics. On the other side, the new degree of freedom of information provided by photon polarization indicates that the $\gamma$-photon polarization could be a useful tool to retrieve the in-situ transient dynamics of plasma electrons. The successful decoding of field properties nearby the event horizon of black holes [90] has re-stimulated the interest in measurements based on photon polarization [91]. While polarized light is vulnerable to magneto-optic disturbance [92], the high-frequency $\gamma$-photon is robust during penetration of the plasma depth [93]. Previously, the celestial $\gamma$-ray emission was observed to understand the star-forming galaxies [94], accretion flows around black holes [95], and active galactic nuclei [96]. In contrast to the routinely detected quantities of arrival time, direction, and energy, the $\gamma$-photon polarization (GPP), provides new insights on the relativistic jet geometry [97] and magnetic field configuration [98], which allows to identify the cosmic neutrino scattering [99], dark matter annihilation [81], and acceleration mechanisms surrounding crab pulsars [100]. All these progresses appeal to a laboratory platform to simulate and examine the GPP associated with the plasma phenomena.

### v2

@@ 3.2 Positrons
The work on polarized positron beams is currently divided between the production and the acceleration of these beams. In contrast to electrons, no pre-polarized targets of positrons can be produced, therefore production via processes like Breit-Wheeler pair creation is necessary. This leaves the realization and verification of the proposed mechanisms up to near-future facilities. Moreover, the efficient acceleration of positrons utilizing wakefields (mostly without considering spin) is still subject of current research, as detailed below.
@@
An alternative method is described by Li et al.: when a high-intensity CP laser pulses interacts with a longitudinally polarized, relativistic electron beam, a highly polarized positron beam can be created via the non-linear Breit-Wheeler process [83]. The polarization is transferred from the electrons to the positrons by high-energy photons. The method is shown to produce positron beams with 40%-60% polarization and low angular divergence on a femtosecond time scale. While this scheme provides a large polarization degree without necessitating high-power two-color laser pulses, a well-polarized electron beam needs to be generated in a first step which is a challenge in itself (cf. previous subsection).
Liu et al. found the polarization of intermediate gamma quanta to be of significance for the polarization of electron-positron pairs produced in the non-linear Breit-Wheeler process [84]. The setup is shown in Fig. 6. They find that an average polarization of 30% can be obtained, and show that the resulting beam can be accelerated in a wakefield driven by a hollow electron beam without significant depolarization.
Figure 6: (a) The hollow driver and an electron seed beam propagate in $z$-direction. (b) The driver beam generates the wakefield structure, and the seed is irradiated by a two-color laser pulse in order to produce positrons. (c) The polarized positrons are then trapped and accelerated in the wakefield. Reproduced under the terms of the $C$-BY license from [84]. Copyright 2022, The Authors, published by American Physical Society.
A method utilizing laser-solid interaction has been proposed by Xue et al. [85]. A high-intensity laser pulse is used to ionize a solid foil. The accelerated electrons are accelerated and generate gamma quanta via non-linear Compton scattering which decay into polarized positrons in a quasi-static magnetic field. The angle of irradiation is found to determine the configuration of the quasi-static field, which in turn, allows tuning of the polarization degree. The scheme delivers polarization on a 70%-level with 0.1 nC per shot. Such solid-target based approaches are simpler to realize as there are no requirements for pre-polarized electron bunches, however, high-intensity laser pulses in the range of $10^{23}\;\mathrm{W/cm}^{2}$ are required, pushing the realization of this setup into near future. Zhao et al. investigated cascaded Compton scattering and Breit-Wheeler pair production on a more general level with respect to polarization transfer in [86].
Another approach to generating polarized positrons was presented by Zhu et al. in [87], using a similar setup to their work on electrons [80]: a relativistic electron beam is used to irradiate a solid target surface, inducing strong, asymmetric magnetic fields. This leads to radiative spin flips and abundant multi-photon processes, generating polarized positrons. In simulations, they find that a multi-GeV positron beam with nC charge and approx. 40% can be produced following this approach. This setup does not require a pre-polarized beam, or high-intensity laser pulses, however, it requires a high-energy electron beam (20 GeV in the case of the presented simulations) that needs to be focused to high density.
The efficient acceleration of positrons using plasma-based methods is – even without considering spin polarization – challenging [88], and still requires significant research efforts for applications like electron-positron colliders [89]. A study by Dou et al. proposed that polarized positrons can be efficiently captured and accelerated from a polarized pair plasma using multilayer microhole-array films [90]. This method preserves polarization on a 90%-level while delivering strong accelerating gradients in the range of TeV/m. While the polarization and energy levels are promising, this method relies on producing a high-quality electron-positron pair plasma first.
@@ 4.1 Magnetic Vortex Acceleration
In the context of polarized protons, MVA was first investigated by Jin et al. in [44] by means of particle-in-cell simulations. They investigated the interaction of a multi-PW laser pulse in the range $a_{0}=25-100$ with a pre-polarized HCl target. The target has an electron density of $0.36n_{\mathrm{crit}}$ and has a flat-top profile with steep edges. In the case of $a_{0}=25$, 0.26 nC of protons with up to 53 MeV and a polarization of 82% were observed. An increase in laser energy was shown to reduce the polarization: for a laser pulse with $a_{0}=100$, 3.1 nC were accelerated to 152 MeV, however with a lower polarization of only 56%.
[…]
@@
In the space between the two plasma channels, an accelerating region with weaker depolarizing fields is formed (cmp. current density in Fig. 7). This leads to the formation of a central filament which obtains energies similar to the comparable single-pulse MVA setup, but with higher polarization. In [97], polarization of 77% was obtained for two pulses with $a_{0}=100$, while a single-pulse setup of the same total laser energy only led to 64%. The dual-pulse scheme was also shown to improve on the angular spectrum of the accelerated beam, as instabilities were mitigated by the presence of the second laser pulse. While the dual-pulse scheme improves on polarization compared to conventional MVA, its experimental realization is more complicated due to the necessity to align to high-power laser pulses sufficiently.
Figure 7: Transverse cut of the current density profile of the dual-pulse MVA scheme. Note the two distinct plasma channels with an accelerating region in their middle. Reproduced under the terms of the $C$-BY license from [97]. Copyright 2022, The Authors, published by American Physical Society.
@@ 4.2 Collisionless Shock Acceleration
[…]
The aforementioned studies have been extended by Reichwein et al. to the case of polarized 3He and to higher laser intensities [101]. A schematic of the setup is shown in Fig. 8. Similarly to the results with protons, high polarization is achieved with CSA, in particular, it remains in the range of 90% for intensities of $a_{0}=$ 100-200. Moreover, the publication investigated the influence of radiation reaction on the acceleration process. As the electrons radiate part of their energy, the electromagnetic field structure of the system and, subsequently, the ion motion is changed. This leads to a reduction in energy, but an increase in beam charge, since more Helium ions can be trapped inside the accelerating field. Still, the beam charge remains in the range of tens of pC, significantly lower than the (up to) few nC that MVA can provide. The polarization remains largely unaffected by radiation reaction.
Figure 8: Schematic of the setup for Collisionless Shock Acceleration. A solid Carbon foil is placed in front of the near-critical density, polarized Helium-3 target. Reproduced under the terms of the $C$-BY license from [101]. Copyright 2024, The Authors, published by IOP Publishing Ltd.
[…]
@@ 4.3 Other mechanisms
[…]
Figure 9: Proton spin polarization for the pure Hydrogen foil (a)-(b), as well as the Aluminium-Hydrogen composite target (c)-(d). Reproduced under the terms of the $C$-BY license from [45]. Copyright 2019, The Authors, published by Cambridge University Press.
@@ 5 Polarized gamma quanta
In this section, we discuss the generation of polarized gamma-ray photons resulting from the interaction of high-intensity laser pulses with plasmas or charged-particle beams. The study of such polarized photons is motivated both by their wide-ranging practical applications and by their potential role in advancing multi-messenger studies of laser-driven plasmas. Polarized gamma-ray photons have practical applications that include nuclear physics [106], high energy physics [107, 108], and astrophysics [109, 110]. Moreover, the photon polarization can provide an additional degree of freedom of information to help retrieve the in-situ electron dynamics in ultrarelativistic laser or beam driven plasmas [111].
@@ 5.1 Polarized Multi-GeV gamma-photon beams via Single-Shot Laser-Electron Interaction
Recently, by considering the photon polarization effect, Li et al. [112] investigated the generation of circularly polarized (CP) and linearly polarized (LP) $\gamma$-rays via the interaction of an ultraintense laser pulse with a spin-polarized counter-propagating ultrarelativistic electron beam. Pre-polarized energetic electrons collide with the laser pulse to experience nonlinear Compton scattering in the quantum radiation dominated regime as shown in Fig. 10. The study develops a Monte Carlo method based on the locally constant field approximation (LCFA) that employs electron-spin-resolved probabilities for polarized photon emissions. Li et al. [112] show efficient ways for the transfer of electron polarization to high-energy photon polarization. The results demonstrate that multi-GeV CP (LP) $\gamma$-rays with polarization of up to about 95% can be generated by a longitudinally (transversely) spin-polarized electron beam, with a photon flux meeting the requirements of recent proposals for the vacuum birefringence measurement in ultra-strong laser fields. The high-energy, high-brilliance, and high-polarization $\gamma$-rays generated in this process are also beneficial for other applications in high-energy particle physics, light sources, and laboratory astrophysics. Recently, using the quantum operator method by Baier and Katkov to calculate the polarization-resolved probabilities within the quasi-classical approach and LCFA, Chen et al. [113] presented a detailed derivation of fully polarization-resolved probabilities for high-energy photon emission and $e^{-}e^{+}$ pair production in ultra-strong laser fields. These probabilities, accounting for both electron spin and photon polarization of incoming and outgoing particles, are crucial for the development of QED Monte Carlo simulations and QED-particle-in-cell codes.
@@
Figure 10: Scenarios of generating CP and LP $\gamma$-rays via nonlinear Compton scattering. (a) An arbitrarily polarized (AP) laser pulse propagating along +z+z direction and head-on colliding with a longitudinally spin-polarized electron bunch produces CP $\gamma$-rays rays. (b) An elliptically polarized laser pulse propagating along +z+z direction and colliding with a transversely spin-polarized (TSP) electron bunch produces LP $\gamma$-rays. The major axis of the polarization ellipse is along the $x$ axis. Reproduced with permission from [112]. Copyright 2020, American Physical Society.
@@
In subsequent studies, Wan et al. [114] investigated the interaction between an unpolarized 10 GeV electron beam and a counter-propagating linearly polarized laser pulse (with a peak intensity $a_{0}\approx 50$ and pulse duration $\tau\approx 50\,$fs) using the same semi-classical Monte Carlo approach. They found that high-energy, linearly polarized $\gamma$-ray photons are generated during the interaction through nonlinear Compton scattering, with an average polarization degree exceeding 50%. These photons further interact with the laser fields, leading to electron-positron pair production via the nonlinear Breit-Wheeler process. Photon polarization is found to enhance the pair production yield by more than 10%. To understand the influence of photon polarization on pair production, Dai et al. [115] used fully polarization-resolved Monte Carlo simulations to examine the correlation between photon and electron (positron) polarization in the multiphoton Breit-Wheeler process, where a 10 GeV electron beam is set to collide with a counter-propagating elliptically polarized laser pulse with a peak intensity $a_{0}=100$, pulse duration $\tau\approx 27\,$fs, and the ellipticity $\epsilon=0.03$. They found that the polarization of the produced $e^{-}e^{+}$ is reduced by 35% and the pair yield is decreased by 13% if the intermediate photon polarization is taken into account.
In addition, Lv et al. proposed to generate brilliant polarized $\gamma$-photon beams via the vacuum dichroism-assisted vacuum birefringence effect, using an unpolarized electron beam [116]. They consider that a linearly polarized laser pulse is split into two sub-pulses, with the first one colliding with a dense unpolarized electron beam to generate a linearly polarized $\gamma$-photon beam, which then further collides with the second sub-pulse and is transformed into a circularly polarized one via the vacuum birefringence effect.
@@ 5.2 Polarized high-energy brilliant gamma-ray sources via laser-plasma interaction
Figure 11: Scenario for the generation of LP $\gamma$-ray photons via nonlinear Compton scattering. An ultra-strong LP laser pulse, polarized along the $y$-axis and propagating along the $x$-direction, irradiates a near-critical density hydrogen plasma followed by an ultra-thin Al foil. Reproduced under the terms of the $C$-BY license from [117]. Copyright 2020, The Authors, published by AIP Publishing.
@@
It is worth emphasizing that the generation of polarized $\gamma$-ray photons via single-shot laser-electron or beam-beam interaction has been studied based on a test-particle simulation [112, 113, 114, 115, 118]. This means that collective effects of charged particles, such as self-generated plasma electric and magnetic fields, are not taken into account. Recently, Xue et al. [117] studied the generation of highly-polarized high-energy brilliant $\gamma$-rays via laser-plasma interaction, where plasma collective effects are self-consistently considered by implementing the electron-spin-resolved probabilities for polarized photon emission in PIC simulations. By utilizing PIC simulations, Xue et al. [117] show that when an ultraintense LP laser pulse irradiates a near-critical-density (NCD) plasma target followed by an ultra-thin aluminum (Al) foil, the electrons in the NCD plasma are first accelerated by the driving laser via direct laser acceleration to ultrarelativistic energies and then collide with the light pulse reflected by the Al foil. The PIC simulations of this all-optical scheme show that brilliant LP $\gamma$-ray photons are produced via nonlinear Compton backscattering with photon energy up to hundreds of MeV and an average LP degree of $\sim 70\%$. This high LP degree is reasonable, since the driving laser pulse is a fully polarized light with a LP degree of $100\%$. Alternatively, instead of using the double-layer target in Fig. 11 to generate $\gamma$-ray photons via nonlinear Compton backscattering, Xue et al. [117] also calculate photon emission by considering a conical gold target filled with an NCD hydrogen plasma. When ultraintense LP light impinges on the NCD target inside the Au cone, the bulk plasma electrons are pushed forward and the consequent current density sustains a strong quasi-static self-generated magnetic field. For forward moving electrons, the self-generated magnetic field provides a restoring force to enable electrons’ betatron oscillation inside a guided plasma channel in the conical Au target, in which emitted $\gamma$-ray photons concentrate in the angular region $\theta<20^{\circ}$, and the cutoff energy exceeds 450 MeV. In this scheme, the LP degree can reach $\sim 50\%$ for the energetic part of emitted photons.
[…]
@@
5.3 Application: deciphering in situ electron dynamics of high-intensity laser-driven plasma
From Li et al. [112] and Xue et al. [117], one can find that $\gamma$-ray photons with high LP degree of $P_{\mathrm{LP}}>70\%$ can be produced via nonlinear Compton backscattering. These highly polarized $\gamma$-ray photons have various applications in laboratory astrophysics and high-energy physics. On the other side, the new degree of freedom of information provided by photon polarization indicates that $\gamma$-photon polarization could be a useful tool for retrieving the in situ transient dynamics of plasma electrons. The successful decoding of field properties near the event horizon of black holes [123] has re-stimulated the interest in measurements based on photon polarization [124]. Although polarized light is vulnerable to magneto-optic disturbance [125], the high-frequency $\gamma$-photon is robust during penetration of the plasma depth [126]. Previously, observation of celestial $\gamma$-ray emission has helped to understand star-forming galaxies [127], accretion flows around black holes [128], and active galactic nuclei [129]. In contrast to routinely detected quantities of arrival time, direction, and energy, the $\gamma$-photon polarization (GPP), provides new insights on the relativistic jet geometry [130] and magnetic field configuration [131], which allows identification of cosmic neutrino scattering [132], dark matter annihilation [110], and acceleration mechanisms surrounding crab pulsars [133]. All this progress appeals to a laboratory platform for simulating and examining the GPP associated with the plasma phenomena.